How do u make a rectangles that have the perimeter of 152
Found 2 solutions by rothauserc, Theo:
Answer by rothauserc(4718)

let l be length and w be width, then
:
2l + 2w = 152
:
divide both sides of = by 2
:
l + w = 76
:
this tells us that the length and width must add up to 76
:
for example,
:
l = 40, w = 36
:
l = 50, w = 26
:
l = 60, w = 16
:
note that l = 38 and w = 38 is also a rectangle because
:
all squares are rectangles but all rectangles are not squares :-)
:

Answer by Theo(13342)

perimeter of a rectangle equals 2L + 2W.
L is the length of the rectangle.
W is the width.
if the perimeter is 152, then 2L + 2W = 152.
factor out the 2 and you get 2 * (L+w) = 152.
divide both sides of this equation by 2 and you get L + W = 152/2 = 76.
solve for L to get L = 76 - W.
the width can be anywhere from 1 to 75 if you are dealing with integer length and widths only.
the area of the rectangle equals L * W.
the perimeter of the rectangle may be the same, but the are will be different depending on the relationship between the length and the width.
the area of the rectangle will be maximum when the rectangle becomes a square because the length and the width is the same.
here are some results from an excel spreadsheet.
you can see that the perimeter is the same but the area is different from this spreadsheet and that the area is greater as the length and the width become closer.
$$$